MsBrains (ms-brains):

PLEASE HELP! BEST RESPONSE AND I'LL HELP YOU BOOST!! Simplify 7x+4y+3(x-y)-2x and Justify your answers! And please show work so I can learn

OpenStudy (anonymous):

=8x+y

OpenStudy (anonymous):

Let's simplify step-by-step. 7x+4y+3(x−y)−2x

OpenStudy (anonymous):

Distribute: =7x+4y+(3)(x)+(3)(−y)+−2x =7x+4y+3x+−3y+−2x

MsBrains (ms-brains):

Yeah, but how did you get that?? I want to learn to do it myself too.. not just the awnser

MsBrains (ms-brains):

@bryan89

OpenStudy (anonymous):

Combine Like Terms: =7x+4y+3x+−3y+−2x =(7x+3x+−2x)+(4y+−3y) =8x+y

OpenStudy (anonymous):

=8x+y

MsBrains (ms-brains):

How do you Justify your answers? @bryan89